a. Gold Is $350 per ounce In the United States and 2.800 pesos per ounce In Mexico. The nominal exchange rate between U.S. dollar and Mexican pesos that is Implled by the PPP theory Is: pesos. b. Mexico experiences Inflation so that the price of gold rises to 4,200 pesos per ounce, whlle the price of gold remalns $350 per ounce In the United States. The nominal exchange rate between U.S. dollars and Mexican pesos that is Implied by the PPP theory is: s-「 pesos. Based on your finding, the following statement is correct: Inflation rates and nominal exchange rates are not correlated. Countries with higher Inflation rates tend to experlence the most rapid appreciation. Countries with higher Inflation rates tend to experience the most rapld depreclation OCountries with lower Inflation rates tend to experlence the most rapld depreclation. c. Assume that gold is $350 per ounce In the United States and 4,200 pesos per ounce In Mexico. Crude oll (excluding taxes and transportation costs) Is $30 per barrel In the Unlted States. According to the PPP theory, one barrel of crude oll should costpesos In Mexico
b. Mexico experiences Inflation so that the price of gold rises to 4,200 pesos per ounce, while the price of gold remains $350 per ounce In the United States. The nominal exchange rate between U.s. dollars and Mexican pesos that is Implied by the PPP theory is: $1-peso. Based on your finding, the following statement Is correct Inflation rates and nominal exchange rates are not correlated. Countries with higher Inflation rates tend to experlence the most rapld appreciation Countrles with higher Inflation rates tend to experlence the most rapld depreclation. OCountries with lower Inflation rates tend to experlence the most rapid depreciation. C. Assume that gold is $350 per ounce In the United States and 4,200 pesos per ounce In Mexico. Crude oll (excluding taxes and transportation costs) IS $30 per barrel In the United States. According to the PPP theory, one barrel of crude oil should cost pesos in Mexico. d. Gold Is $350 per ounce in the United States. The exchange rate between the United States and Canada Is 0.70 U.S. dollars per Canadlan dollar. Therefore, one ounce of gold should cost in Canadian dollars
Answer #1

a) nominal exchange rate= 2800/350 = 8

Therefore, $1=8 pesos

b) new nominal exchange rate= 4200/350= 12

Therefore, $1= 12 pesos

Based on the findings the following statement is correct: Countries with higher inflation rates ted to experience the most rapid depreciation.

c) As we know, $1= 12 pesos

Now, Price of crude oil in US= $30 per barrel

Hence, Price of crude oil in Mexico= 30*12= 360 pesos

d) Price of gold in US= $350 per ounce

Exchange rate between US and Canada = 0.70

Thus, price of gold in Canada (per ounce) = 350/0.70= $500